Multichoice Ghana offers two basic packages: Sports and Movies. There are three

different types of users: parents, sports fans and generalists. The table below shows

the maximum price each type of consumer is willing to pay for each package.

Consumer type Sports (GHC) Movies (GHC)

Parents 10 40

Sports fans 50 10

Generalists 40 40

a) What price should Multichoice Ghana charge for each package and what will be

its total revenue if it cannot bundle the two packages?

b) If Multichoice wants to bundle the two packages, what price should it charge?

What will be its total revenue?

c) If Multichoice decides to sell each package for GHC 50 and a combined bundle

for GHC 70. How does the revenue from this method compare with those in parts

(a) and (b)?


1
Expert's answer
2022-03-11T08:32:35-0500

a)

Sports package

"10+50+40=GHC\space 100"


Movies package

"40+10+40=GHC\space 90"


Total revenue will be;

"100+ 90=GHC\space190"


b)

If Multichoice wants to bundle the two packages it should charge GHC 50. This is because the three types of users will purchase them since they will be within their budget lines. Parents are willing to spend GHC 50, Sports fans GHC 60 while Generalists are willing to spend GHC 80.


Total revenue will be;

"50+50+50= GHC\space 150"


c)

If Multichoice decides to sell each package for GHC 50, Only sports fans will buy the sports package because it is within their budget and they will get the most use out of it. Consumers want items that are within their budget and that they can use often.

Both packages are desired by generalists, who are willing to pay $40 for each. Both bundles are available for $70 when purchased together. The generalists will pay $70 for both packages because it is less than their full budget and they will get exactly what they want.


Total revenue

"50+70= GHC\space 120"


GHC 120 will be the total revenue which is lesser compared to that of (a) and (b)